OpenStudy (anonymous):

A=2(pi)r^2+2(pi)hr, for r

OpenStudy (anonymous):

\[A=2pir ^{2}+2pihr \]

OpenStudy (perl):

its not easy to solve for r

OpenStudy (perl):

you can use quadratic formula

OpenStudy (perl):

where did you get this problem

OpenStudy (anonymous):

my math homework on web assign

OpenStudy (perl):

i think you have a mistake. what was the original question

OpenStudy (perl):

yeah thats a bit odd

OpenStudy (perl):

@ganeshie8 can you take a look at this

OpenStudy (perl):

can you post the other parts of the question

OpenStudy (anonymous):

that is all.

OpenStudy (perl):

could be a typo

OpenStudy (perl):

@dan815 can you look at this

OpenStudy (perl):

@erikuh i will get ask around, im sure somebody has an idea

OpenStudy (perl):

@rvc can you check this out

rvc (rvc):

gtg @perl

OpenStudy (anonymous):

okay thankyou(: i will ask my teacher if its a typo

OpenStudy (dan815):

ya why dont u use quadratic formula

OpenStudy (dan815):

A=2PIR^2+2PIHR a=kr^2+khr 0=kr^2+khr-a \[r= \frac{-kh +/- sqrt( k^2h^2-4*k*-a)}{2k}\]